# A comprehensive comparison of neural operators for 3D industry-scale engineering designs

### v1

@@ Abstract
Neural operators have emerged as powerful tools for learning nonlinear mappings between function spaces, enabling real-time prediction of complex dynamics in diverse scientific and engineering applications. With their growing adoption in engineering design evaluation, a wide range of neural operator architectures have been proposed for various problem settings. However, model selection remains challenging due to the absence of fair and comprehensive comparisons. To address this, we propose and standardize six representative 3D industry-scale engineering design datasets spanning thermal analysis, linear elasticity, elasto-plasticity, time-dependent plastic problems, and computational fluid dynamics. All datasets include fully preprocessed inputs and outputs for model training, making them directly usable across diverse neural operator architectures. Using these datasets, we conduct a systematic comparison of four types of neural operator variants, including Branch-Trunk-based Neural Operators inspired by DeepONet, Graph-based Neural Operators inspired by Graph Neural Networks, Grid-based Neural Operators inspired by Fourier Neural Operators, and Point-based Neural Operators inspired by PointNet. We further introduce practical enhancements to adapt these models to different engineering settings, improving the fairness of the comparison. Our benchmarking study evaluates each model’s strengths and limitations in terms of predictive performance, computational efficiency, memory usage, and deployment complexity. The findings provide actionable insights to guide future neural operator development. All source code is available to the public at Github. The data is available at Harvard Dataverse for the paper reviewing process: Heat sink, Bracket, Bracket-time, JEB, DrivAer, and DrivAer++.
@@ 3 Evaluation framework on industry-scale engineering problems
A neural operator learns a mapping between infinite-dimensional function spaces. Let $D\subset\mathbb{R}^{d}$ be a Lipschitz domain, and let $\mathcal{B}$ be a Banach space of real-valued functions defined on $D$. Consider a family of level set functions $\mathcal{S}\subset\mathcal{B}$, where each $s\in\mathcal{S}$ defines a $d$-dimensional submanifold $\Omega_{S}=\{x\in D:s(x)>0\}$. Given a partial differential operator $\mathcal{L}$, we study the solution $u\in\mathcal{U}_{T}\subset\mathcal{B}$ of the following PDE:
@@
  • Bracket-time: A time-dependent extension of the bracket dataset, where varying loading sequences are applied over multiple time steps. The goal is to learn the transient stress evolution in elasto-plastic materials.
  • JEB: The Jet Engine Bracket (JEB) dataset consists of 2,138 geometrically diverse 3D bracket designs sourced from the GE Jet Engine Bracket Challenge. The shapes exhibit significant freeform variation without a shared parameterization. The task is to predict the von Mises stress field under a prescribed vertical load.
  • DrivAerNet: A large-scale CFD dataset offering high-fidelity aerodynamic simulations across multiple 3D vehicle geometries with parametric variation. The objective is to predict surface pressure distribution under standardized wind flow conditions.
  • DrivAerNet++: An extension of DrivAerNet involving freeform, non-parametric vehicle geometries. This dataset focuses on evaluating model generalization to unparameterized, complex CAD geometries in automotive fluid dynamics.
@@ 4.1 Branch-Trunk-Based Neural Operators
where $\mathcal{U}_{b}$ is a multi-layer perceptron (MLP) applied to the input parameter $p$, $\mathcal{U}_{t}$ is an MLP applied to the spatial query $x_{q}$, and $\odot$ represents the element-wise multiplications. However, the simplicity of this architecture limits its effectiveness in addressing complex engineering problems. Consequently, several extensions have been proposed to enhance the performance of the original DeepONet and to adapt it to a broader range of problem types.
For model performance improvement, several variants of DeepONet have been proposed. Geom-DeepONet replaces the branch MLP with a sinusoidal representation network (SIREN), showing the higher model capacity to encode intricate geometric features. The Deep Compositional Operator Network (DCON) increases model expressiveness by introducing a hierarchical composition of operator layers. For example, with two layers, the solution is approximated by:
@@
where $\mathcal{U}_{u}^{1}$ and $\mathcal{U}_{u}^{2}$ are learnable MLPs. This compositional structure improves representation capacity and has been shown to achieve higher accuracy on complex PDEs.
As for problems with time-dependent PDEs, S-DeepONet replaces the branch MLP with a gated recurrent unit (GRU), enabling temporal sequence encoding in the branch net:
@@
where $f(t_{k})$ denotes a time-dependent scalar function values at time step $t_{k}$. This temporal modeling makes S-DeepONet particularly effective in transient simulations, such as dynamic load responses in structural systems. A more advanced version using attention mechanism is also proposed for time series problem Liu and Koric (2025).
To enable handling of free-form geometries, Geometry-aware Neural Operator (GANO) extends DCON by incorporating point-cloud-based geometric encoding, where a global feature vector $E_{g}$ is concatenated with the trunk input:
@@
where $\parallel$ denotes concatenation, and $\mathcal{U}_{g}$ encodes each geometry point $x_{i}$. GANO improves accuracy on non-parametric domains but introduces preprocessing overhead for geometry encoding, which can be challenging to integrate into legacy engineering workflows.
Overall, these branch-trunk models provide complementary strengths for different engineering problems. DeepONet and DCON are suitable for structured parametric problems, Geom-DeepONet shows strong capability in handling parametric geometry variability, S-DeepONet supports temporal dynamics, and GANO is adapted to free-form geometries. All these models are selected as the representative of branch-trunk models for engineering design evaluation.
@@ 4.2 Graph-Based Neural Operators
Graph Neural Operators (GNOs) approximate mappings between function spaces using message passing over graphs, making them suitable for solving PDEs on irregular meshes. A typical GNO predicts the PDE solution on the mesh point coordinate $x_{i}$ with stacked graph convolutions:
@@
where $\mathcal{N}_{v}$ is the set of neighbor nodes, $\mathcal{U}_{\text{mp}}$ is the message passing function, and $\mathcal{U}_{f}$ encodes the neighbor features. Edge-Augmented GNNs (EA-GNNs) improve long-range information flow by introducing edges between distant node pairs. These are often heuristically or randomly added in point clouds, reducing graph diameter. From a deployment perspective, both models are practical for engineering applications, as they can directly leverage existing meshes generated from finite element analysis without requiring additional pre-processing.
Graph Transformer Operators substitute traditional convolution with attention mechanisms to better capture long-range dependencies, computing attention scores across multi-hop neighbors, often based on $k$-hop shortest path distances. Although this enhances the ability to model global correlations, the computational cost of shortest path calculations makes them impractical for integration into large-scale PDE solvers. Similarly, Multipole Graph Neural Operators leverage hierarchical graph representations to reduce complexity, but they rely on constructing multi-resolution graphs through pairwise distance computations for each sample, which imposes significant overhead and hinders real-time deployment. In this study, we only investigate GNOs and EA-GNOs, the graph-based models that can operate directly on FEM meshes easily, without remeshing or new graph construction, ensuring minimal disruption to engineering pipelines.
@@ 4.3 Grid-Based Neural Operators
Fourier Neural Operator (FNO) initiates the grid-based operator learning approach by performing convolution in the spectral domain, which predicts the PDE solution on a grid by:
@@
where $G^{l}$ is the grid feature output of the $l$-th Fourier layer, $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively, and $R_{\phi}$ is a learned spectral filter. Similarly, F-FNO factorizes the spectral weights for improved efficiency, and U-NO adopts a U-Net structure for hierarchical multiscale features. However, all these grid-based models are restricted to uniform discretizations, which severely limits its application to general engineering problems involving curved boundaries or unstructured meshes. To overcome this, Geo-FNO maps irregular domains into a latent uniform grid through a learnable transformation, enabling FFT-based processing However, it still struggles to generalize across drastically different shapes.
Geometry-Informed Neural Operator (GINO) extends FNO by integrating point cloud information directly. It uses a Riemann-sum-based discretization to construct the mapping between unstructured point features and grid-aligned features:
$v_{g}$ $=\sum_{\|x_{g}-x_{i}\|_{2}\leq r}\mathcal{U}_{\text{mp}}(v^{0}_{i},v_{g})\,\mathcal{U}_{f}(v^{0}_{i}),\quad v_{g}\in G^{0},$ (14)
$G^{l}$ $=\mathcal{U}^{l}_{\text{FL}}\circ\dots\circ\mathcal{U}^{1}_{\text{FL}}(G^{0}),$ (15)
$v^{l}_{i}$ $=\sum_{\|x_{g}-x_{i}\|_{2}\leq r}\mathcal{U}_{\text{mp}}(v_{g},v^{l}_{i})\,\mathcal{U}_{f}(v_{g}),\quad v_{g}\in G^{l},$ (16)
where $r$ is the hyper-parameter to search the neighbor nodes of each grid point, $v_{i}$ and $v_{g}$ represent node and grid features, respectively. However, storing and processing a high-resolution 3D grid is computationally expensive and memory inefficient. To alleviate this, FigConv-UNet uses factorized implicit grids instead of explicit 3D grids and processes data with a U-Net backbone. This drastically improves memory efficiency and has demonstrated strong performance on large CFD meshes.
From a deployment perspective, although the method requires radius search operations, these can be efficiently executed using hash grid algorithms available in GPU-accelerated libraries such as Warp. Therefore, we consider that both GINO and FigConv-UNet are suitable for large-scale engineering applications.
@@ 4.4 Point-Based Neural Operators
Point-based neural operators aim to directly model solution operators on point clouds. They offer high flexibility and are naturally suited to domains with nonuniform resolution and geometry. PointNet processes unordered sets using shared MLPs and symmetric pooling, which is easily deployed but lacks local structure modeling.
PointNet++ improves upon PointNet by incorporating local neighborhoods for hierarchical feature extraction. While these models are simple and computationally efficient, they require careful point cloud grouping and sampling, which complicates their integration into general-purpose engineering design workflows. Similarly, PCNO and DoMINO enhance performance on unstructured domains by introducing local geometry-aware modules and multi-scale decomposition strategies. However, these improvements come with significant preprocessing overhead, making them less practical for deployment within conventional CAD-to-solver pipelines.
More recently, transformer-based architectures have emerged, predicting PDE solution for one point location with stacked attention layers:
@@
where different models propose different attention mechanism. For example, GNOT introduces global linear attention to capture nonlocal dependencies:
@@
where $\tau$ is the hyper-parameter, $Q^{l}_{i}=\mathcal{U}_{q}(v^{l}_{i})$, $K^{l}_{j}=\mathcal{U}_{k}(v^{l}_{j})$, and $V^{l}_{j}=\mathcal{U}_{v}(v^{l}_{j})$ represents the queries, keys, and values of the point features. To further enhance the efficiency of transformer nerual operators, Transolver proposes a linear self-attention mechanism using physics-aware slice tokens. Each point feature $v_{j}$ is assigned to $M$ slices using a MLP $\mathcal{U}_{\text{slice}}$:

### v2

@@ Abstract
Neural operators have emerged as powerful tools for learning nonlinear mappings between function spaces, enabling real-time prediction of complex dynamics in diverse scientific and engineering applications. With their growing adoption in engineering design evaluation, a wide range of neural operator architectures have been proposed for various problem settings. However, model selection remains challenging due to the absence of fair and comprehensive comparisons. To address this, we propose and standardize six representative 3D industry-scale engineering design datasets spanning thermal analysis, linear elasticity, elasto-plasticity, time-dependent plastic problems, and computational fluid dynamics. All datasets include fully preprocessed inputs and outputs for model training, making them directly usable across diverse neural operator architectures. Using these datasets, we conduct a systematic comparison of four types of neural operator variants, including Branch-Trunk-based Neural Operators inspired by DeepONet, Graph-based Neural Operators inspired by Graph Neural Networks, Grid-based Neural Operators inspired by Fourier Neural Operators, and Point-based Neural Operators inspired by PointNet. We further introduce practical enhancements to adapt these models to different engineering settings, improving the fairness of the comparison. Our benchmarking study evaluates each model’s strengths and limitations in terms of predictive performance, computational efficiency, memory usage, and deployment complexity. The findings provide actionable insights to guide future neural operator development. All source code is available to the public at Github. The data is available at Harvard Dataverse for the paper reviewing process: Heat sink, Bracket, Bracket-time, JEB, and DrivAer and DriverNet++
@@ 3 Evaluation framework on industry-scale engineering problems
A neural operator learns a mapping between infinite-dimensional function spaces. Let $D\subset\mathbb{R}^{d}$ be a Lipschitz domain Cobzaş et al. (2019), and let $\mathcal{B}$ be a Banach space of real-valued functions defined on $D$. Consider a family of level set functions Osher and Fedkiw (2001) $\mathcal{S}\subset\mathcal{B}$, where each $s\in\mathcal{S}$ defines a $d$-dimensional submanifold $\Omega_{S}=\{x\in D:s(x)>0\}$. Given a partial differential operator $\mathcal{L}$, we study the solution $u\in\mathcal{U}_{T}\subset\mathcal{B}$ of the following PDE:
@@
  • Bracket-time: A time-dependent extension of the bracket dataset, where varying loading sequences are applied over 101 time steps. The goal is to learn the final stress state of the bracket in elasto-plastic modeling.
  • JEB: The Jet Engine Bracket (JEB) dataset consists of 2,138 geometrically diverse 3D bracket designs sourced from the GE Jet Engine Bracket Challenge Hong et al. (2025). The shapes exhibit significant freeform variation without a shared parameterization. The task is to predict the von Mises stress field under a prescribed vertical load.
  • DrivAer: A large-scale CFD dataset offering high-fidelity aerodynamic simulations across multiple 3D vehicle geometries with parametric variation Elrefaie et al. (2024a). The objective is to predict surface pressure distribution under standardized wind flow conditions.
  • DrivAer++: An extension of DrivAerNet involving freeform, non-parametric vehicle geometries Elrefaie et al. (2024b). This dataset focuses on evaluating model generalization to unparameterized, complex CAD geometries in automotive fluid dynamics.
@@ 4.1 Branch-Trunk-Based Neural Operators
where $\mathcal{U}_{b}$ is a multi-layer perceptron (MLP) applied to the input parameter $p$, $\mathcal{U}_{t}$ is an MLP applied to the spatial query $x_{q}$, and $\odot$ represents the element-wise multiplications. However, the simplicity of this architecture limits its effectiveness in addressing complex engineering problems. For model performance improvement, Geom-DeepONet He et al. (2024a) replaces the branch MLP with a sinusoidal representation network (SIREN). On the other side, the Deep Compositional Operator Network (DCON) Zhong and Meidani (2024) increases model expressiveness by introducing a hierarchical composition of operator layers, For example, with two layers, the solution is approximated by:
@@
where $\mathcal{U}_{u}^{1}$ and $\mathcal{U}_{u}^{2}$ are learnable MLPs. As for problems with time-dependent PDEs, S-DeepONet He et al. (2024b) replaces the branch MLP with a gated recurrent unit (GRU), enabling temporal sequence encoding in the branch net:
@@
where $f(t_{k})$ denotes a time-dependent scalar function values at time step $t_{k}$. To enable handling of free-form geometries, Geometry-aware Neural Operator (GANO) Zhong and Meidani (2025) extends DCON by incorporating point-cloud-based geometric encoding, where a global feature vector $E_{g}$ is concatenated with the trunk input:
@@
where $\parallel$ denotes concatenation, and $\mathcal{U}_{g}$ encodes each geometry point $x_{i}$. Overall, compared with traditional DeepONet, these improved branch-trunk models provide complementary strengths for different engineering problems.
@@ 4.2 Graph-Based Neural Operators
Graph Neural Operators (GNOs) Li et al. (2020c) approximate mappings between function spaces using message passing over graphs, making them suitable for solving PDEs on irregular meshes. A typical GNO predicts the PDE solution on the mesh point coordinate $x_{i}$ with stacked graph convolutions:
@@
where $\mathcal{N}_{v}$ is the set of neighbor nodes, $\mathcal{U}_{\text{mp}}$ is the message passing function, and $\mathcal{U}_{f}$ encodes the neighbor features. Edge-Augmented GNNs (EA-GNNs) Gladstone et al. (2024) improve long-range information flow by introducing edges between distant node pairs. These are often heuristically or randomly added in point clouds, reducing graph diameter. From a deployment perspective, both models are practical for engineering applications, as they can directly leverage existing meshes generated from finite element analysis without requiring additional pre-processing.
@@ 4.3 Grid-Based Neural Operators
Fourier Neural Operator (FNO) Li et al. (2020b) initiates the grid-based operator learning approach by performing convolution in the spectral domain, which predicts the PDE solution on a grid by:
@@
where $G^{l}$ is the grid feature output of the $l$-th Fourier layer, $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively, and $R_{\phi}$ is a learned spectral filter. Geometry-Informed Neural Operator (GINO) Li et al. (2023c) extends FNO by integrating point cloud information directly. It uses a Riemann-sum-based discretization to construct the mapping between unstructured point features and grid-aligned features:
$v^{1}_{g}$ $=\sum_{\|x_{g}-x_{i}\|_{2}\leq r}\mathcal{U}_{\text{mp}}(v^{0}_{i},v^{0}_{g})\,\mathcal{U}_{f}(v^{0}_{i}),\quad v^{1}_{g}\in G^{1},\quad v^{0}_{g}\in G^{0},\quad v^{0}_{i}\in X,$ (14)
$G^{l}$ $=\mathcal{U}^{l}_{\text{FL}}\circ\dots\circ\mathcal{U}^{1}_{\text{FL}}(G^{1}),$ (15)
$v^{l}_{i}$ $=\sum_{\|x_{g}-x_{i}\|_{2}\leq r}\mathcal{U}_{\text{mp}}(v^{l}_{g},v^{0}_{i})\,\mathcal{U}_{f}(v^{l}_{g}),\quad v^{l}_{i}\in X^{l},\quad v^{l}_{g}\in G^{l},$ (16)
where $r$ is the hyper-parameter to search the neighbor nodes of each grid point, $v_{i}$ and $v_{g}$ represent node and grid features, respectively. However, storing and processing a high-resolution 3D grid is computationally expensive and memory inefficient. To alleviate this, FigConvNet Choy et al. (2025) uses factorized implicit grids instead of explicit 3D grids and processes data with a U-Net backbone, which drastically improves memory efficiency and has demonstrated strong performance on large CFD meshes. From a deployment perspective, although the method requires radius search operations, these can be efficiently executed using hash grid algorithms Sahebi et al. (2025) available in GPU-accelerated libraries such as Warp. Therefore, we consider that both GINO and FigConvNet are suitable for large-scale engineering applications.
@@ 4.4 Point-Based Neural Operators
Point-based neural operators aim to directly model solution operators on point clouds. They offer high flexibility and are naturally suited to domains with nonuniform resolution and geometry. PointNet Kashefi and Mukerji (2022) processes unordered sets using shared MLPs and symmetric pooling, which is easily deployed but lacks local structure modeling. More recently, transformer-based architectures Hao et al. (2023) have emerged, predicting PDE solution for one point location with stacked attention layers:
@@
where the attention mechanism is formulated as:
@@
where $\tau$ is the hyper-parameter, $Q^{l}_{i}=\mathcal{U}_{q}(v^{l}_{i})$, $K^{l}_{j}=\mathcal{U}_{k}(v^{l}_{j})$, and $V^{l}_{j}=\mathcal{U}_{v}(v^{l}_{j})$ represents the queries, keys, and values of the point features. To further enhance the efficiency of transformer nerual operators, Transolver Wu et al. (2024) proposes a linear self-attention mechanism using physics-aware slice tokens. Each point feature $v_{j}$ is assigned to $M$ slices using a MLP $\mathcal{U}_{\text{slice}}$:
@@ 7 Conclusion
Acknowledgment
This work used the Delta advanced computing resources provided by the University of Illinois Urbana-Champaign (UIUC) through the National Center for Supercomputing Applications (NCSA). We gratefully acknowledge the support of these resources.